d^2y/dx^2 + 3(dy/dx) + 2y = sin(e^x )
1
Expert's answer
2020-06-16T15:53:37-0400

"k^2+3k+2=0"

"k_1=-2"

"k_2=-1"

"y(x)=C_1e^{k_1x}+C_2e^{k_2x}"

"y(x)=C_1e^{-2x}+C_2e^{-x}"


"e^{-x}\frac {d}{dx}C_2(x)+e^{-2x}\frac {d}{dx}C_1(x)=0"

"\frac {d}{dx}(e^{-x})\frac {d}{dx}C_2(x)+\frac {d}{dx}(e^{-2x})\frac {d}{dx}C_1(x)=sin(e^x)"


"e^{-x}\frac {d}{dx}C_2(x)+e^{-2x}\frac {d}{dx}C_1(x)=0"

"-e^{-x}\frac {d}{dx}C_2(x)-2e^{-2x}\frac {d}{dx}C_1(x)=sin(e^x)"


"\frac {d}{dx}C_1(x)=-e^{2x}sin(e^x)"

"\frac {d}{dx}C_2(x)=e^{x}sin(e^x)"


"C_1(x)=-\int e^{2x}sin(e^x)dx=e^xcos(e^x)-sin(e^x)+C_3"

"C_2(x)=\int e^{x}sin(e^x)dx=C_4-cos(e^x)"


Answer:

"y(x)=C_3e^{-2x}+C_4e^{-x}-e^{-2x}sin(e^x)"